### v2

@@ 1 Introduction
In recent years, machine learning has evolved as an appropriate solution for fake news detection for several reasons. Firstly, for a machine learning model to work, it needs to learn with a significant amount of data. And with the proliferation of social media and online news sources, there is now a significant amount of data available to train machine learning models to detect fake news. Machine learning algorithms process large amounts of data automatically and find underlying patterns or characteristics which can help detect fake news. Pattern recognition is particularly useful as many fake news share an underlying pattern. Research in fake news has been conducted for a long time, especially in English. There has been extensive work on classifying fake news in the English language bib20 ; bib21 ; bib22 ; bib23 ; bib24 ; bib25 ; bib26 . Recently other than the English language, research on fake news is being conducted in Spanish bib27 , French bib28 ; bib29 ; bib30 ; bib31 , Arabic bib32 , Chinese bib33 ; bib34 ; bib35 ; bib36 ; bib37 ; bib38 ; bib39 , and many other languages bib40 ; bib41 ; bib42 ; bib43 ; bib44 ; bib45 ; bib46 ; bib47 ; bib48 ; bib49 . There are many different methods used in these researches but deep learning based methods seemed to be the most popular.
@@ 2.1 Traditional Approaches
Using traditional machine learning based approach, there are several techniques to detect fake news. This approach typically involves feature engineering and the use of various classifiers such as Support Vector Machines (SVM), Logistic Regression (LR), Naive Bayes (NB), Random Forests (RF), K-nearest Neighbor algorithm (KNN), Stochastic Gradient Descent (SGD), Passive Aggressive Classifier (PAC) and Decision Trees (DT).
@@
Khan et al. bib61 conducted a benchmark study to evaluate the effectiveness of several practical approaches on three distinct English datasets, the largest and most diverse of which was created by them. They experiment with both traditional machine learning based and neural network based approaches. Within traditional models they used SVM, LR, DT, MNB and KNN - and among them Naive Bayes, with n-gram features, performed the best.
Baarir et al. bib62 used two existing English datasets and divided news data into three categories - Textual, categorical and numerical. Category wise, they used different preprocessing techniques. They only used Support Vector Machine (SVM) classifier but experimented with different hyper-parameters to improve accuracy. Another unique aspect of their research is that, they experimented with different features and through experimentations denote that the best features are in this order - text, author, source, date and sentiment. Using all features together they achieved 100% accuracy in their research.
@@
Hossain et al. (2020) bib10 created BanFakeNews, which is an annotated dataset of around 48 thousand authentic and around 1300 fake news in Bengali. To our knowledge, this is the only publicly available dataset in on Bengali fake news and it paved research on Bengali fake news detection. They collected authentic news articles from 22 reputed sources and three types of fake news - misleading, clickbait and satire are collected from some popular websites. They initially experiment with traditional machine learning models - SVM, LR and RF and empirically find that SVM incorporated with all linguistic features outperforms the other two models, achieving 0.91 F1-score on fake class. They also observe that lexical features perform better than other linguistic features. Additionally, they claim that the use of punctuation in fake news is more frequent, and most of the time fake news is found on the least popular sites.
@@
Hussain et al. (2020) bib63 employed Multinomial Naive Bayes (MNB) and Support Vector Machine (SVM) algorithms to construct a Bengali fake news detection model. The model leveraged both count vectorizer and TF-IDF for feature extraction. They created their own dataset of around 2500 news articles among which 1548 were authentic and 993 fake. In their experimentation, SVM with an accuracy of 0.9664 outperforms MNB with an accuracy of 0.9332.
@@
Sraboni et al. (2021) bib64 employed the BanFakeNews dataset bib10 alongside an additional dataset of 2,500 news articles bib63 for training various traditional machine learning algorithms, including Random Forest (RF), Passive-Aggressive Classifier (PAC), Multinomial Naive Bayes (MNB), Support Vector Machine (SVM), Logistic Regression (LR), and Decision Tree (DT). TF-IDF served as the feature extraction method. Rather than using all of 51.8k available data, they used 3.5k authentic data and 2.3k fake data to remove bias. They also experimented with different train-test split ratios with 50-50, 60-40, 80-20 and 70-30 and empirically found that 70-30 split ratio gives the best result. Among the models, Passive Aggressive Classifier (PAC) and Support Vector Machine (SVM) achieve 0.938 and 0.935 accuracy respectively which are higher than the other models they trained.
@@
Mugdha et al. bib68 created their own balanced dataset of total 112 instances with an equal number of fake and real news. They tested multiple traditional algorithms with tenfold cross validation - Support Vector Machine (SVM), Logistic Regression (LR), Multi-layer Perceptron (MLP), Random Forest Classifier (RF), Voting Ensemble Classifier (VEC) and Gaussian Naive Bayes (GNB). Initially, none of the algorithms were performing well. Then they applied 3 feature selection techniques - PCA, Kernel PCA and Extra Tree Classifier. Through empirical research, they discovered that Extra Tree Classifier outperformed other techniques and that GNB with Extra Tree Classifier outperformed other models, with an accuracy of 0.8552.
@@
Piya et al. (2022) bib72 investigated the development of a model capable of concurrent identification of Bengali and English fake news. Their approach for detecting fake news from a bilingual perspective uses feature extraction techniques such as TF-IDF and N-Gram analysis. They trained six supervised traditional machine learning models - Logistic Regression, Linear SVC, Decision Tree Classifier, Random Forest Classifier, Multinomial Naive Bayes, and Passive Aggressive Classifier. Among these six models, the linear SVC performed best and score 0.9329 in accuracy and 0.93 F1-score. They adopted BanFakeNews for the Bengali portion of their Dataset. Additionally, they used ISOT, a dataset released in bib66 , for the English dataset. In their final dataset, they had 42,324 bilingual samples. The majority of them were in English.
@@
Mughal et al. (2020) bib73 concentrated on leveraging headlines alone for Bengali fake news detection. They developed their own Bengali dataset with a total of 538 cases, of which 269 were authentic and 269 were false. They have used 9 different traditional machine learning algorithms for their research - SVM, Logistic Regression(LR), multilayer perceptron (MLP), Random Forest Classifier (RF), Voting Ensemble Classifier (VEC), Gaussian Naive Bayes (GNB), Multinomial Naive Bayes (MNB), AdaBoost (AB) and Gradient Boosting (GB). With a score of 0.87, the Gaussian Naive Bayes classifier outperformed all other models in terms of accuracy. This algorithm chose the attribute using an Extra Tree Classifier and a text feature dependent on TF-IDF. Performance was assessed using stratified 10-fold cross-validation on the dataset. They contribute to this field by creating a novel dataset and also a new Bengali stemmer.
@@ 2.2 Neural Network based Approaches
Sharma et al. bib11 adopted Convoluted Neural Network (CNN) based approach built upon a hybrid feature extraction model. They created their own dataset by crawling satire news portal - Motikontho for satire news and real news from Prothom ALo and Ittefaq, two popular news media. They gather 1480 satire news and to balance the dataset, they picked the same amount from the real news. With this balanced training, they were able to achieve 0.96 in accuracy.
Bahad et al. bib75 proposed using Bi-directional LSTM recurrent neural network to detect fake news. They used two publicly available news articles datasets separately. They experimented with Bi-directional LSTM, CNN, vanilla RNN, unidirectional LSTM and found Bi-directional LSTM to be superior, achieving .9108 in accuracy.
In the paper of Khan et al. bib61 , along with traditional methods, they also experimented with deep learning models such as - CNN, LSTM, Bi-LSTM, C-LSTM, HAN, Conv-HAN, char level C-LSTM. Their interpretation of the result is that no deep learning model is uniquely superior to others as no model performs consistently well on all three of their datasets. In addition, they note that while neural network-based models exhibit good accuracy and f1-score on a fairly large dataset (Combined Corpus), they exhibit overfitting on a small dataset (LIAR). Keya et al. in their paper - AugFakeBERT bib12 experimented with all three - traditional, deep learning and transformer based approach. For deep learning models they used CNN, LSTM, Bi-LSTM and CNN-LSTM.
@@ 2.3 Hybrid and Ensemble Approaches
Ruchansky et al. bib69 propose a hybrid model named CSI consisting of three modules - Capture, Score and Integrate. The first module employs a recurrent neural network (RNN) to capture the temporal structure of user activity on a particular article and is based on the response and text. The second module picks up on the source characteristics based on user activity, and then the two are combined with the third module to determine whether or not an article is authentic. They used two datasets - TWITTER and WEIBO. Their hybrid CSI model was able to achieve 0.95 in accuracy and F1-score.
The research team of Nasir et al. bib70 proposed a novel hybrid deep learning model for classifying fake news that integrates recurrent and convolutional neural networks. They used two publicly available datasets FA-KES and ISOT. Both datasets have almost equal amounts of fake and real news. They were able to achieve an accuracy score of 0.60 in dataset-1 and 0.99 in dataset-2.
The research of Ahmad et al. bib66 proposed ensembling multiple traditional machine learning models. They picked three publicly available English datasets which contain news from multiple domains. They explored various ensemble techniques such as bagging, boosting and voting classifier. Their first voting classifier is ensemble of LR,RF and KNN and second one consists of LR, linear SVM, and CART. The bagging ensemble consists 100 decision trees and two boosting algorithms are XGBoost and AdaBoost. Comparing ensemble and individual learners, the ensemble learners have consistently outperformed the individual learners in their research.
@@ 2.4 Transformer based Approaches
Saha et al. bib67 used two Bengali dataset and experimented with Pre-trained BERT transformer, LSTM with regularization, CNN, SVM and NB. They used two Bengali datasets separately for training and testing. In their research, BERT achieves 94% F1-score which is superior to other methods explored.
@@ 3 Corpus Description
In this section, we provide a brief description of the corpora that we used for training and test purpose. Throughout our research, we used a total of three corpora: two publicly available corpora and a manually collected corpus which we refer to as - ‘CustomFake’.
(a) BanFakeNews Corpus: BanFakeNews bib10 dataset contains approximately 50,000 Bengali news articles among which about 48 thousand are of authentic category and 1299 are of fake category. The dataset contains seven attributes - articleID, domain, date, category, headline, content and label. The label here is binary where 0 denotes fake class and 1 denotes authentic class. For our research, we merged the headlines and their corresponding news content. We show few statistics of the corpus in Table 2.
Table 2: BanFakeNews dataset statistics
Authentic Fake
Number of instances 48678 1299
Average length of articles 1858.95 1817.73
Average word count 283.6 289.54
Longest article word count 4791 2713
BERT Max token length 12539 8107
[…]
@@ 4.3 Augmentation
    Back translation: Back translation method works as follows: we take some sentences (e.g. in Bengali) and translate to another language (e.g English). Then we translate the English sentences back to Bengali sentences. The final Bengali sentences will have subtle changes than the original Bengali sentences. We used the publicly available bnaug [^9] https://github.com/sagorbrur/bnaug library to perform back translation.
@@ 5.3 Hyper-parameters Tuning
Hyper-parameters are parameters that are set before the training process begins such as the learning rate, batch size, loss function, activation function, the number of epochs etc. The hyper-parameters are adjusted until the best performance is achieved on the validation set. We report the hyper-parameter settings for all our experiments in Table 8. As our train and test datasets are balanced, we consider accuracy and f1-score as the primary metrics for assessing the superiority of the models. Since all classes have an equal number of samples, we specifically focus on the macro average. Moreover, we take note of other scores such as precision, recall, MCC, and ROC-AUC to gain a more comprehensive understanding of the performance of each model.

### v3

@@ 1 Introduction
In recent years, machine learning has evolved as an appropriate solution for fake news detection for several reasons. Firstly, for a machine learning (ML) model to work, it needs to learn with a significant amount of data. And with the proliferation of social media and online news sources, there is now a significant amount of data available to train machine learning models to detect fake news. Machine learning algorithms process large amounts of data automatically and find underlying patterns or characteristics which can help detect fake news. Pattern recognition is particularly useful as many fake news share an underlying pattern. Research in fake news has been conducted for a long time, especially in English. There has been extensive work on classifying fake news in the English language bib20 ; bib21 ; bib22 ; bib23 ; bib24 ; bib25 ; bib26 . Recently other than the English language, research on fake news is being conducted in Spanish bib27 , French bib28 ; bib29 ; bib30 ; bib31 , Arabic bib32 , Chinese bib33 ; bib34 ; bib35 ; bib36 ; bib37 ; bib38 ; bib39 , and many other languages bib40 ; bib41 ; bib42 ; bib43 ; bib44 ; bib45 ; bib46 ; bib47 ; bib48 ; bib49 . There are many different methods used in these researches but deep learning based methods seemed to be the most popular.
@@
To assess the performance of the proposed approaches, we address the following research questions which are discussed in detail in section 5.4 of our paper:
  1. (a)
  2. (b)
  3. (c)
  4. (d)
@@ 2.1 Traditional Approaches
Using traditional machine learning (ML) based approach, there are several techniques to detect fake news. This approach typically involves feature engineering and the use of various classifiers such as Support Vector Machines (SVM), Logistic Regression (LR), Naive Bayes (NB), Random Forests (RF), K-nearest Neighbor algorithm (KNN), Stochastic Gradient Descent (SGD), Passive Aggressive Classifier (PAC) and Decision Trees (DT).
@@
Khan et al. bib61 conducted a benchmark study to evaluate the effectiveness of several practical approaches on three distinct English datasets, the largest and most diverse of which was created by them. They experiment with both traditional ML based and neural network based approaches. Within traditional models they used SVM, LR, DT, Multinomial Naive Bayes (MNB) and KNN - and among them Naive Bayes, with n-gram features, performed the best.
Baarir et al. bib62 used two existing English datasets and divided news data into three categories: Textual, categorical and numerical. They employed an SVM classifier and experimented with various feature extraction and hyperparameter configurations to enhance accuracy. They identified the most impactful features in the following order: text, author, source, date, and sentiment, ultimately achieved 100% accuracy when using all features together. However, such a result is typically considered improbable in real-world classification tasks and may indicate presence of overfitting, where the model memorizes the training data rather than learning generalizable patterns. Their use of source and author features for model training and evaluation may introduce bias. For example, if a dataset predominantly associates certain sources, such as 100percentfedup.com, with fake news, the model may learn to classify news based on the source rather than its content, thereby reducing generalizability.
@@
Hossain et al. bib10 created BanFakeNews, which is an annotated dataset of around 48 thousand authentic and around 1300 fake news in Bengali. To our knowledge, this is the only publicly available dataset in on Bengali fake news and it paved research on Bengali fake news detection. They collected authentic news articles from 22 reputed sources and three types of fake news - misleading, clickbait and satire are collected from some popular websites. They initially experiment with traditional ML models - SVM, LR and RF and empirically find that SVM incorporated with all linguistic features outperforms the other two models, achieving 0.91 $\ignorespaces F_{1}$ -score on fake class. They also observe that lexical features perform better than other linguistic features. Additionally, they claim that the use of punctuation in fake news is more frequent, and most of the time fake news is found on the least popular sites.
@@
Hussain et al. bib63 employed MNB and SVM algorithms to construct a Bengali fake news detection model. The model leveraged both count vectorizer and TF-IDF for feature extraction. They created their own dataset of around 2500 news articles among which 1548 were authentic and 993 fake. In their experimentation, SVM with an accuracy of 0.9664 outperforms MNB with an accuracy of 0.9332. While the results are promising, the relatively small size of the dataset raises concerns about the model’s ability to generalize effectively to real-world scenarios involving larger and more diverse texts.
@@
Sraboni et al. bib64 employed the BanFakeNews dataset bib10 alongside an additional dataset of 2,500 news articles bib63 for training various traditional ML algorithms, including RF, PAC, MNB, SVM, LR, and DT. TF-IDF served as the feature extraction method. Rather than using all of 51.8k available data, they used 3.5k authentic data and 2.3k fake data to remove bias. They also experimented with different train-test split ratios with 50-50, 60-40, 80-20 and 70-30 and empirically found that 70-30 split ratio gives the best result. Among the models, PAC and SVM achieve 0.938 and 0.935 accuracy respectively which are higher than the other models they trained. However, TF-IDF is a non-contextual method and does not capture the sequential or semantic relationships between words, limiting the model’s deeper understanding of the textual content.
@@
Mugdha et al. bib68 created their own balanced dataset of total 112 instances with an equal number of fake and real news. They tested multiple traditional algorithms with tenfold cross validation - SVM, LR, Multi-layer Perceptron (MLP), RF, Voting Ensemble Classifier (VEC) and Gaussian Naive Bayes (GNB). Initially, none of the algorithms were performing well. Then they applied 3 feature selection techniques - Principal Component Analysis (PCA), Kernel PCA and Extra Tree Classifier (ETC). Through empirical research, they discovered that ETC outperformed other techniques and that GNB with ETC outperformed other models, with an accuracy of 0.8552. While Mugdha et al. explored various traditional ML algorithms and feature selection techniques, their dataset was relatively small, consisting of only 112 instances.
@@
Piya et al. bib72 investigated the development of a model capable of concurrent identification of Bengali and English fake news. Their approach for detecting fake news from a bilingual perspective uses feature extraction techniques such as TF-IDF and N-Gram analysis. They trained six supervised traditional ML models - LR, Linear SVC, DT, RF, MNB, and PAC. Among these six models, the linear SVC performed best and score 0.9329 in accuracy and 0.93 $\ignorespaces F_{1}$ -score. They adopted BanFakeNews for the Bengali portion of their Dataset. Additionally, they used ISOT, a dataset released in bib66 , for the English dataset. In their final dataset, they had 42,324 bilingual samples. The majority of them were in English.
@@
Mugdha et al. bib73 concentrated on leveraging headlines alone for Bengali fake news detection. They developed their own Bengali dataset with a total of 538 cases, of which 269 were authentic and 269 were false. They have used 9 different traditional ML algorithms for their research - SVM, LR, MLP, RF, VEC, GNB, MNB, AdaBoost (AB) and Gradient Boosting (GB). With a score of 0.87, the GNB classifier outperformed all other models in terms of accuracy. This algorithm chose the attribute using an ETC feature extractor and a text feature dependent on TF-IDF. Performance was assessed using stratified 10-fold cross-validation on the dataset. They contribute to this field by creating a novel dataset and also a new Bengali stemmer. However, the research is limited by the modest size of the training dataset and its singular focus on headline-based analysis.
@@ 2.2 Neural Network based Approaches
Sharma et al. bib11 adopted Convolutional Neural Network (CNN) based approach built upon a hybrid feature extraction model. They created their own dataset by crawling satire news portal - Motikontho for satire news and real news from Prothom ALo and Ittefaq, two popular news media. They gather 1480 satire news and to balance the dataset, they picked the same amount from the real news. For vectorization, They used 1000-dimensional TF-IDF vectorizer and 10-dimensional Word2Vec which resulted in 1000 x 10 x 2 (2D embeddings with separate positive and negative channels) document representation. With this balanced training, they were able to achieve 0.96 in accuracy.
Bahad et al. bib75 proposed using Bi-directional LSTM recurrent neural network to detect fake news. They used two publicly available news articles datasets separately. They also experimented with CNN, vanilla Recurrent Neural Network (RNN), unidirectional LSTM and found Bi-LSTM to be superior, achieving 0.9108 in accuracy.
In the paper of Khan et al. bib61 , along with traditional methods, they also experimented with deep learning models such as - CNN, LSTM, Bi-LSTM, Convolutional-LSTM (C-LSTM), Hierarchical Attention Network (HAN), Conv-HAN, char-LSTM. For the CNN, LSTM, Bi-LSTM models, 100 dimensional pre-trained GloVe embedings were used and HAN and Conv-HAN architectures employed 100 dimensional bidirectional GRU encoding. Their interpretation of the result is that no deep learning model is uniquely superior to others as no model performs consistently well on all three of their datasets. In addition, they note that while neural network-based models exhibit good accuracy and $\ignorespaces F_{1}$ -score on a fairly large dataset (Combined Corpus), they exhibit overfitting on a small dataset (LIAR). Keya et al. in their paper - AugFakeBERT bib12 experimented with all three - traditional, deep learning and transformer based approach. For deep learning models they used CNN, LSTM, Bi-LSTM and CNN-LSTM with 128 dimensional trainable embeddings.
@@ 2.3 Hybrid and Ensemble Approaches
Ruchansky et al. bib69 propose a hybrid model named CSI, consisting of three modules—Capture, Score, and Integrate—and utilize two datasets: TWITTER and WEIBO. The first module employs an RNN to capture the temporal structure of user activity on a particular article, based on both user responses and textual content. User features ($x_{u}$) are embedded using Singular Value Decomposition (SVD), with ranks of 20 for Twitter and 10 for Weibo. The second module focuses on source characteristics derived from user activity, and the outputs of the first two modules are combined in the third module to determine the authenticity of an article. For Weibo, textual features ($x_{\tau}$) are extracted using doc2vec applied to segmented text. The initial feature vector $x_{t}$ has a dimensionality of 122 for Twitter and 112 for Weibo. These raw features are projected into a lower-dimensional embedding space of 100 dimensions ($\tilde{x}_{t}$) before being passed into the RNN. Their hybrid CSI model was able to achieve 0.95 in accuracy and $\ignorespaces F_{1}$ -score.
The research team of Nasir et al. bib70 proposed a novel hybrid deep learning model for classifying fake news that integrates recurrent and convolutional neural networks. They used two publicly available datasets FA-KES and ISOT and employed 100 dimensional pre-trained GloVe embeddings. Both datasets have almost equal amounts of fake and real news. They were able to achieve an accuracy score of 0.60 in dataset-1 and 0.99 in dataset-2.
The research of Ahmad et al. bib66 proposed ensembling multiple traditional ML models. They picked three publicly available English datasets which contain news from multiple domains. They explored various ensemble techniques such as bagging, boosting and voting classifier. Their first voting classifier is ensemble of LR,RF and KNN and second one consists of LR, linear SVM, and Classification and Regression Trees (CART). The bagging ensemble consists 100 DT and two boosting algorithms are XGBoost and AB. Comparing ensemble and individual learners, the ensemble learners have consistently outperformed the individual learners in their research.
@@ 2.4 Transformer based Approaches
Saha et al. bib67 used two Bengali dataset and experimented with Pre-trained BERT transformer, LSTM with regularization, CNN, SVM and NB. They used two Bengali datasets separately for training and testing. In their research, BERT achieves 94% $\ignorespaces F_{1}$ -score which is superior to other methods explored.
A significant limitation in previous work lies in the common reliance on TF-IDF for feature extraction, which fails to capture contextual relationships within the text. Moreover, the limited availability of large and balanced fake news datasets often leads to the development of models susceptible to biases due to imbalanced training data. To overcome these critical shortcomings, our research aims to create a more comprehensive and balanced dataset through collection, translation, and augmentation, and to leverage the contextual understanding capabilities of the BERT model for improved fake news detection.
@@ 3 Corpus Description
In this section, we provide a brief description of the corpora that we used for training and test purposes. Throughout our research, we used a total of three corpora: two publicly available corpora and a manually collected corpus which we refer to as - ‘CustomFake’.
(a) BanFakeNews Corpus: BanFakeNews bib10 dataset contains approximately 50,000 Bengali news articles among which about 48 thousand are of authentic category and 1299 are of fake category. The dataset contains seven attributes - articleID, domain, date, category, headline, content and label. The label here is binary where 0 denotes fake class and 1 denotes authentic class. For our research, we merged the headlines and their corresponding news content. We show a few statistics of the corpus in Figure 1.
Figure 1: Pipeline of the summarization process.
@@ 4.3 Augmentation
    Back translation: Back translation method works as follows: we take some sentences (e.g. in Bengali) and translate to another language (e.g English). Then we translate the English sentences back to Bengali sentences. The final Bengali sentences will have subtle changes than the original Bengali sentences. We used the publicly available bnaug [^9] https://github.com/sagorbrur/bnaug library to perform back translation. This technique preserves the semantic meaning of the original text while introducing syntactic and lexical variations. It is particularly effective for low-resource languages like Bengali, where other augmentation methods may struggle to maintain linguistic integrity pingle2023robust .
@@ 5.1 Evaluation Metrics
 $\text{ recall }=\frac{TP}{TP+FN}$ (3)
  • $\ignorespaces F_{1}$
Score: The $\ignorespaces F_{1}$ Score is the harmonic mean of precision and recall. In our binary classification task, we evaluate model performance by computing the macro average of metrics such as the $\ignorespaces F_{1}$ score. The macro average is the arithmetic mean of the $\ignorespaces F_{1}$ scores calculated independently for each class.
  • F1
=2precisionrecallprecision+recall
@@ 5.3 Hyper-parameters Tuning
Hyper-parameters are parameters that are set before the training process begins such as the learning rate, batch size, loss function, activation function, the number of epochs etc. The hyperparameters were fine-tuned to achieve the best performance on the validation set. Table 7 presents the final hyperparameter settings used across all experiments, while Figure 5 illustrates the relationship between training and validation loss versus accuracy. As our train and test datasets are balanced, we consider accuracy and $\ignorespaces F_{1}$ -score as the primary metrics for assessing the superiority of the models. Since all classes have an equal number of samples, we specifically focus on the macro average. Moreover, we take note of other scores such as precision, recall, MCC, and ROC-AUC to gain a more comprehensive understanding of the performance of each model.
Figure 5: Training and validation loss vs Accuracy for TM-mBERT.